OpenStudy (anonymous):

9 root 7 difference root 28

OpenStudy (anonymous):

\[9\sqrt{7} -\sqrt{28}\] you can simplify this to \[9\sqrt{7} - \sqrt{4 \times 7}\] which becomes \[9\sqrt{7} - 2\sqrt{7}\] let y = sqrt(7) so now its a simple subtraction than substitute back. \[9y - 2y = ?\]

OpenStudy (anonymous):

7y?

OpenStudy (anonymous):

good now substitute back y

OpenStudy (anonymous):

7 root7

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

thank you and love your pic btw

OpenStudy (anonymous):

yay XD your welcome =))